Please help me solve this equation
(2x+1)(3x-2) = 55

Answer by ikdeep(189)

(2x+1)(3x-2) = 55..
here you must know how to solve...if you might not know then please contact me again ,,and i will explain to you)...
solving of the brackets would give us
6x^2 -4x +3x - 2 = 55
6x^2 -x - 2 = 55
or 6x^2 -x - 2 - 55 = 0
(Here you must know how we make factors i.e. splitting of the middle term...if you might not know then please contact me again and I will explain to you... )
6x^2 -x - 57 = 0
6x^2 -(19-18)x - 57 = 0

6x^2 -19x + 18x - 57 = 0
taking common from first two i.e.6x^2 -19x and last two i.e. 18x - 57 variables...we get
x^2(6x-19) - 3 (6x-19) = 0
or (x^2-3) (6x-19) = 0
hope this will help ...
Please feel free to revert back for any further queries.